What is the percent by mass of water in iron(II) s
Answer #1

Molar mass of FeSO4.7H2O = 55.845 + 32 + 4 * 16 + 7 * (16+1+1)

=> 55.845 + 32 + 64 + 126

=> 277.845 gm/mol

Mass of water in FeSO4.7H2O = 7 * 18 = 126 gm/mol

Percentage of water as mass = 126/277.845 * 100 = 45.34%